OpenStudy (anonymous):

MEDAL AS A REWARD!!! The area of a regular hexagon is 55 in.2 Find the length of a side. Round your answer to the nearest tenth. I need help figuring out how to solve this!

OpenStudy (dumbcow):

use area formula for regular hexagon http://en.wikipedia.org/wiki/Hexagon#Regular_hexagon

OpenStudy (anonymous):

.................I still don't get it?

OpenStudy (dumbcow):

\[55 = \frac{3\sqrt{3}}{2}s^{2}\]

OpenStudy (anonymous):

.......

OpenStudy (anonymous):

I'm seriously stuck :C

OpenStudy (dumbcow):

are you having trouble using the formula? or are you trying to understand how to derive the formula

OpenStudy (anonymous):

both :'C

OpenStudy (dumbcow):

ok  hexagon has 6 equal triangles...find area of 1 triangle  from trig, tan 60 = sqrt3 tan = opp/adj = h/(x/2) = sqrt3 --> h = (sqrt3/2)x Area of 1 triangle = 1/2*x*h = (sqrt3/4)x^2 multiply by 6 --> Area = (3sqrt3/2)x^2

OpenStudy (anonymous):

So I'm basically working my way backwards.....okay